40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 16 03:43:09 2014, Jenne, Update, LSC, PRMI loops need help PRMI_55vs165_15Dec2014.pdf
    Reply  Tue Dec 16 20:51:18 2014, diego, Update, LSC, PRMI loops need help 
    Reply  Wed Dec 17 13:14:38 2014, ericq, Update, LSC, PRMI loops need help oops.pdf
       Reply  Wed Dec 17 14:42:13 2014, Jenne, Update, LSC, PRMI loops need help 
Message ID: 10804     Entry time: Tue Dec 16 03:43:09 2014     Reply to this: 10806   10809
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI loops need help 

[Jenne, Rana, Diego]

After deciding that the Yend QPD situation was not significant enough to prevent us from locking tonight, we got started.  However, the PRMI would not acquire lock with the arms held off resonance. 

This started some PRMI investigations.

With no arms, we can lock the PRMI with both REFL55 I&Q or REFL165 I&Q.  We checked the demod phase for both Refl 55 and 165.  REFL55 did not need changing, but REFL165 was off significantly (which probably contributed to the difficulty in using it to acquire lock).  I didn't write down what REFL165 was, but it is now -3 degrees.  To set the phase (this is also how Rana checked the 55 phase), I put in an oscillation using the sensing matrix oscillators.  For both REFL165I and 165Q, I set the sensing matrix demod phases such that all of the signal was in the I phase (so I_I and Q_I, and basically zero in I_Q and Q_Q).  Then, I set the main PD demod phase so that the REFL165Q phase (the Q_I phase) was about zero.

Here are the recipes for PRMI-only, REFL55 and REFL165:

Both cases, actuation was PRCL = 1*PRM and MICH = (0.5*BS - 0.2625*PRM).  Trigger thresholds for DoFs and FMs were always POP22I, 10 up and 0.5 down.

REFL55, demod phase = 31deg.

MICH = 2*R55Q, gain = 2.4, trig FMs 2, 6, 8.

PRCL = 12*R55I, gain = -0.022, trig FMs 2,6,9.

REFL165, demod phase = -3deg.

MICH = -1*R165Q, gain = 2.4, trig FMs 2,6,8.

PRCL = 2.2*R165I, gain = -0.022, trig FMs 2,6,9.

These recipes assume Rana's new resonant gain filter for MICH's FM6, with only 2 resonant gains at 16 and 24 Hz instead of a whole mess of them: elog 10803.  Also, we have turned down the waiting time between the MICH loop locking, and the filters coming on.  It used to be a 5 second delay, but now is 2 sec.  We have been using various delays for the PRCL filters, between 0.2s and 0.7s, with no particular preference in the end.

We compared the PRCL loop with both PDs, and note that the REFL 165 error signal has slightly more phase lag, although we do not yet know why.  This means that if we only have a marginally stable PRCL loop for REFL55, we will not be stable with REFL165. Also, both loops have a non-1/f shape at a few hundred Hz.  This bump is still there even if all filters except the acquisition ones (FM4,5 for both MICH and PRCL) are turned off, and all of the violin filters are turned off.  I will try to model this to see where it comes from.

PRMI_55vs165_15Dec2014.pdf

To Do list:

Go back to the QPDY situation during the daytime, to see if tapping various parts of the board makes the noise worse.  Since it goes up to such high frequencies, it might not be just acoustic.  Also, it's got to be in something common like the power or something, since we see the same spectra in all 4 quadrants. 

The ASS needs to be re-tuned. 

Rana was talking about perhaps opening up the ETMX chamber and wiggling the optic around in the wire.  Apparently it's not too unusual for the wire to get a bit twisted underneath, which creates a set of places that the optic likes to go to.

Diego is going to give us some spectra of the MC error point at various levels of pockel's cell drive.  Is it always the same frequencies that are popping up, or is it random?

ELOG V3.1.3-